# Data Mixing Laws: Optimizing Data Mixtures by Predicting Language Modeling Performance

### v1

@@ 5 Application to Continual Pretraining Outlooks Data Schedule Driven by Data Mixing Laws
During continual pretraining, a too-large proportion of the target data can hurt the performance of the original data. A representative mixture optimization target is to maintain the general-purpose ability (looses on the Pile) unchanged. To this end, using the fitted data mixing laws, we predict the critical proportion leading to the same loss as before continual pretraining. Fig. 8 demonstrates the success of our prediction where the proportion we find results in similar performance compared to the model before continual pretraining while gaining improvement in the target domain.
@@ 6 Related Work
Training massive transformer architecture on trillions of tokens, a.k.a. pretraining, is the primary step to building modern large language models that exhibit impressive human-like generalist abilities (Brown et al., 2020; OpenAI, 2023; Jiang et al., 2023; Touvron et al., 2023b)). It takes up most of the computation resources for model training and researchers believe it endows almost all the knowledge in LLMs (Zhou et al., 2024). Such critical impact motivates the development of data curating strategies to reduce computation costs and enhance knowledge (Longpre et al., 2023). The efforts can be categorized into two steps. The first step focuses on obtaining a high-quality training dataset. A typical procedure includes selecting data sources to constitute different domains, deduplication, and the most intricate filtering (Wenzek et al., 2019; Penedo et al., 2023). A mass of endeavors in existing practice has involved multifarious filters, scoring the documents with from superficial features on characters (Rae et al., 2021; Xie et al., 2024b; Raffel et al., 2020) to semantics including similarity to the high-quality reference corpus (Wenzek et al., 2019) and toxicity (Longpre et al., 2023; Friedl, 2023). With a dataset on hold, the second step aims to make the best use of it. This includes tuning the data mixture (Du et al., 2022; Touvron et al., 2023a; Xie et al., 2024a) and devising data schedules (Mindermann et al., 2022; Albalak et al., 2023; Chen et al., 2024b; Fan et al., 2024). Our work is among those tune data mixtures and our extension to continue pretraining signifies our prospect of guiding the schedule design. Different from existing attempts that rely on intuition or qualitative targets, our study seeks a quantitative solution.
@@ 7 Discussions
In this work, we discover the quantitative predictability of model losses regarding the mixture proportions of training data, which boils down to the data mixing laws. Our findings cover training data from two domains to multiple domains, and validation data from a single domain to unknown compositions. Using data mixing laws allows practitioners to quantitatively estimate the model performance on unseen mixture proportions before the actual training, facilitating the tuning of ideal mixture proportions. We further propose the nested use of scaling laws of training steps, model sizes, and data mixture to make predictions with only experiments at small scales, which enables the reuse of existing experiments and reduces the computation costs. Experimental results suggest that our method effectively optimizes the data mixture, which leads to better performance in pretraining and maintaining the original abilities of pretrained models in continual pretraining.
Admittedly, how data mixtures affect model training is far from fully understood. Our study makes preliminary attempts at a quantitative framework while leaving several limitations.
@@
The concept of domains is not well-defined. In this paper, similar to related studies (Xie et al., 2024a; Chen et al., 2024b; Albalak et al., 2023; Fan et al., 2024), we directly adopt the predefined domains in the open-source training data. Nevertheless, we suppose that more operationally defined training domains, e.g., clustering (Gururangan et al., 2023; Shao et al., 2024), could further benefit the prediction accuracy of data mixing laws and the performance of outcome models. For the validation domains, our implicit domain aggregation method obviates the necessity of explicitly aligning validation data with training domains. This requirement is often encountered, given that validation data typically comprises trustworthy datasets rather than mere compilations from training domains. However, we acknowledge that implicit domain aggregation may be less interpretable compared to the explicit approach and may raise concerns regarding its accuracy, as elaborated subsequently.
@@
To sum up, we initiate an investigation into quantitative prediction methods for pretraining data curation. With its increasing focus on data engineering, we hope our exploration facilitates further quantitative studies and theoretical analyses in this research area.
Acknowledgments
Appendix A The relative qualities of data mixtures relate to model sizes and training steps.
@@ Acknowledgments
Results show that the relative performance fluctuates despite a relatively consistent trend across sizes and training steps. This indicates that a mixture is better at small scales probably but does not always perform better at large scales. The longest common sequence of the partial orders among the 20 mixtures in Fig. 9(A) and Fig. 9(B) only reaches lengths of 10 and 11, respectively.
Appendix B Implementation Details
B.1 Model Training
Throughout this study, we employ the Pythia Suit (Biderman et al., 2023) as our model architectures, the specific configurations are in Tab. 2. The maximum sequence length is 4096 for pretraining from scratch and 2048 for continual pretraining, where the latter aligns with the setting of the original pretrained models. In all our experiments, we train the model with a batch size of 1M tokens and a maximum learning rate of 1e-4. We warm up the learning rates for 2000 steps and decay it to 0.1 of the maximum at the last training step with a cosine decay schedule. For continual pretraining, we initialize the models with the 20k-step checkpoint of the Pythia 70M model and do not apply a learning rate warmup.
@@ B.1 Model Training
B.2 Fitting Mixture Laws
@@
Appendix C Connections between Implicit Domain Aggregation and MLP
@@
Inspired by this perspective, we attribute the successful fitting of mixture law to two aspects. First, the MLP with a sufficiently large hidden dimension is a universal approximator (Hornik et al., 1989) thus being able to fit the relationships between losses and mixture proportions. Second, the mixture proportions are bounded between 0 and 1. For this reason, predicting an unseen mixture is an interpolation problem, which is usually easier than extrapolation.

### v2

@@ 5 Application to Continual Pretraining Outlooks Data Schedule Driven by Data Mixing Laws
During continual pretraining, a too-large proportion of the target data can hurt the performance of the original data. A representative mixture optimization target is to maintain the general-purpose ability (losses on the Pile) unchanged. To this end, using the fitted data mixing laws, we predict the critical proportion leading to the same loss as before continual pretraining. Fig. 10 demonstrates the success of our prediction where the proportion we find results in similar performance compared to the model before continual pretraining while gaining improvement in the target domain.
@@ 6 Related Work
Training massive transformer architecture on trillions of tokens, a.k.a. pretraining, is the primary step to building modern large language models that exhibit impressive human-like generalist abilities (Brown et al., 2020; OpenAI, 2023; Jiang et al., 2023; Touvron et al., 2023b; Sun et al., 2024)). It takes up most of the computation resources for model training and researchers believe it endows almost all the knowledge in LLMs (Zhou et al., 2024). Such critical impact motivates the development of data curating strategies to reduce computation costs and enhance knowledge (Longpre et al., 2023). The efforts can be categorized into two steps. The first step focuses on obtaining a high-quality training dataset. A typical procedure includes selecting data sources to constitute different domains, deduplication, and the most intricate filtering (Wenzek et al., 2019; Penedo et al., 2023). A mass of endeavors in existing practice has involved multifarious filters, scoring the documents with from superficial features on characters (Rae et al., 2021; Xie et al., 2024b; Raffel et al., 2020) to semantics including similarity to the high-quality reference corpus (Wenzek et al., 2019) and toxicity (Longpre et al., 2023; Friedl, 2023). With a dataset on hold, the second step aims to make the best use of it. This includes tuning the data mixture (Du et al., 2022; Touvron et al., 2023a; Xie et al., 2024a) and devising data schedules (Mindermann et al., 2022; Albalak et al., 2023; Chen et al., 2024b; Fan et al., 2024). Our work is among those tune data mixtures and our extension to continue pretraining signifies our prospect of guiding the schedule design. Different from existing attempts that rely on intuition or qualitative targets, our study seeks a quantitative solution. As a concurrent work, Liu et al. (2024) also proposes to predict optimal data mixtures through regression but assuming rank invariance across training scales.
@@ 7 Discussions
In this work, we discover the quantitative predictability of model losses regarding the mixture proportions of training data, which boils down to the data mixing laws. Using data mixing laws allows practitioners to quantitatively estimate the model performance on unseen mixture proportions before the actual training, allowing low-cost tuning of data mixture together with scaling laws. Given the burgeoning interest in data engineering, we hope that our study paves the way for further quantitative inquiries and theoretical analyses in this research area.
Acknowledgments
Appendix A Related Work
Curating pretraining data for LLMs.
Training massive transformer architecture on trillions of tokens, a.k.a. pretraining, is the primary step to building modern large language models that exhibit impressive human-like generalist abilities (Brown et al., 2020; OpenAI, 2023; Jiang et al., 2023; Touvron et al., 2023b; Sun et al., 2024)). It takes up most of the computation resources for model training and researchers believe it endows almost all the knowledge in LLMs (Zhou et al., 2024). Such critical impact motivates the development of data curating strategies to reduce computation costs and enhance knowledge (Longpre et al., 2023). The efforts can be categorized into two steps. The first step focuses on obtaining a high-quality training dataset. A typical procedure includes selecting data sources to constitute different domains, deduplication, and the most intricate filtering (Wenzek et al., 2019; Penedo et al., 2023). A mass of endeavors in existing practice has involved multifarious filters, scoring the documents with from superficial features on characters (Rae et al., 2021; Xie et al., 2024b; Raffel et al., 2020) to semantics including similarity to the high-quality reference corpus (Wenzek et al., 2019) and toxicity (Longpre et al., 2023; Friedl, 2023). With a dataset on hold, the second step aims to make the best use of it. This includes tuning the data mixture (Du et al., 2022; Touvron et al., 2023a; Xie et al., 2024a) and devising data schedules (Mindermann et al., 2022; Albalak et al., 2023; Chen et al., 2024b; Fan et al., 2024). Our work is among those tune data mixtures and our extension to continue pretraining signifies our prospect of guiding the schedule design. Different from existing attempts that rely on intuition or qualitative targets, our study seeks a quantitative solution. As a concurrent work, Liu et al. (2024) also proposes to predict optimal data mixtures through regression but assuming rank invariance across training scales.
Scaling laws
Appendix B Limitations and Discussions
How data mixtures affect model training is far from fully understood. Our study makes preliminary attempts at a quantitative framework while leaving several limitations.
@@ Appendix B Limitations and Discussions
The concept of domains is not well-defined. In this paper, similar to related studies (Xie et al., 2024a; Chen et al., 2024b; Albalak et al., 2023; Fan et al., 2024), we directly adopt the predefined domains in the open-source training data. Nevertheless, we suppose that more operationally defined training domains, e.g., clustering (Gururangan et al., 2023; Shao et al., 2024), could further benefit the performance of outcome models. For the validation domains, our implicit domain aggregation method obviates the necessity of explicitly aligning validation data with training domains. This requirement is often encountered, given that validation data typically comprises trustworthy datasets rather than mere compilations from training domains. However, we acknowledge that implicit domain aggregation may be less interpretable compared to the explicit approach and may raise concerns regarding its accuracy, as elaborated subsequently.
@@
We believe that further investigation into these issues could facilitate more sophisticated quantitative methods for data engineering. We leave them as future work.
Appendix C The ranking of data mixtures depend on model sizes and training steps.
@@
Results show that the relative performance fluctuates despite a relatively consistent trend across sizes and training steps. This indicates that a mixture is better at small scales but does not always perform better at large scales, consistent with findings of Goyal et al. (2024); Covert et al. ; Kang et al. (2024). The longest common sequence of the partial orders among the 20 mixtures in Fig. 11(A) and Fig. 11(B) only reaches lengths of 10 and 11, respectively.
Appendix D Implementation Details
D.1 Model Training
Throughout this study, we employ the Pythia Suit (Biderman et al., 2023) as our model architectures, the specific configurations are in Tab. 2. The maximum sequence length is 4096 for pretraining from scratch and 2048 for continual pretraining, where the latter aligns with the setting of the original pretrained models. In all our experiments, we train the model with a batch size of 1M tokens and a maximum learning rate of 1e-4. We warm up the learning rates for 2000 steps and decay it to 0.1 of the maximum at the last training step with a cosine decay schedule. For continual pretraining, we initialize the models with the 20k-step checkpoint of the Pythia 70M model and do not apply a learning rate warmup. For the costs of our experiments, it takes around 3.5/8/16/21 hours to train a 70M/160M/305M/410M model for 30B tokens on 8 A100 GPUs on our infrastructure.
@@ D.1 Model Training
D.2 Predicting Language Modeling Performance with Scaling Laws
In our prediction pipeline in Sec. 4, we adopt nested use scaling laws of training steps and model sizes, which are both power laws, to predict language modeling performance at scale. To fit the laws, we follow Hoffmann et al. (2022) to search over a set of initialized parameters and fit the samples by minimizing the Huber errors between predictions and observations with LBFGS.
We present our results on verifying the feasibility of applying scaling laws to predict language modeling performance. Our prediction pipeline (described in Sec. 4) employs two scaling laws: one related to training steps and another to model sizes, to extrapolate performance with increased training data and larger models. We evaluate the precision of predictions for each of these scaling laws, respectively.
Scaling laws of training steps.
Fig. 12 shows the training curve 70M models on three different data mixtures. We fit power laws within 30k steps (marked with circles) and extrapolate to predict model performance to as many as 100k steps (marked with stars). On all validation sets, the fitted curves give descent prediction precision, with a low mean absolute error of 0.02.
Figure 12: Verification on predicting language modeling performance with scaling laws of training steps. We train 70M models on three different mixtures up to 100k steps and validate them on 5 validation domains as well as the overall validation mixture. All curves are fitted within 30k steps (•) and and extrapolated to predict model performance to 100k steps ($\star$)
Scaling laws of model sizes.
Fig. 13 shows the results where we fit power laws on 70M, 160M, and 305M models (marked with circles) and extrapolate the curve to predict 410M model performance (marked with stars) at different training steps and under different data mixtures. The results are positive, showing that we can precisely predict the 410M model performance at different training steps, with a mean absolute error of 0.003.
Figure 13: Verification on predicting language modeling performance with scaling laws of model sizes. We train models of 70M, 160M, and 406M on three different mixtures and validate them on 5 validation domains as well as the overall validation mixture. All curves are fitted with models of 70M, 160M, and 305M (•) and extrapolated to predict the performance of 410M models ($\star$). We verify the predictability at different numbers of training steps.
Overall, we consider fitting power laws to predict model performance for more training steps and larger models are feasible. Therefore we adopt them to implement the scaling law predictions in our pipeline (Sec. 4).
D.3 Fitting Data Mixing Laws
@@
Appendix E Connections between Implicit Domain Aggregation and MLP
@@
Inspired by this perspective, we attribute the successful fitting of data mixing laws to two aspects. First, the MLP with a sufficiently large hidden dimension is a universal approximator (Hornik et al., 1989) thus being able to fit the relationships between losses and mixture proportions. Second, the mixture proportions are bounded between 0 and 1. For this reason, predicting an unseen mixture is an interpolation problem, which is usually easier than extrapolation.
@@
Appendix F Supplemented Results
F.1 Prediction Results on More Domains
To further consolidate the efficacy of data mixing laws and show that they are general for different data, we experiments on domains different from those in Sec. 3.2.
We train and validate on Wikipedia, ArXiv, and StackExchange of RedPajama, which are three domains different from those in Sec. 3.2. All samples are from 70M models trained for 10k steps. The prediction accuracy is in Fig. 15. The result shows the predicted and observed losses are consistent for different mixtures. This confirms that our data mixing laws also work on domains besides those in the main experiments.
Figure 15: Prediction results on domain losses with Eqn. 7. We train 70M models on mixtures of Wikipedia, ArXiv, and StackExchange for 10k steps. We fit on 7 mixtures and validate on 3 other mixtures.
F.2 Data Mixing Laws Make no Domain-Independent Assumptions
Although our data mixing laws combine the terms with the proportion of different domains through linear combination, we make no domain-independent assumption that different domain affects the losses independently. This is because the linear combination serves as an exponent in Eqn. 6 and Eqn. 7. Specifically, by Taylor expansion, we have
$L_{i}(r_{1\dots M})=c_{i}+k_{i}\exp{\left(\sum_{j=1}^{M}{t_{ij}r_{j}}\right)}=c_{i}+k_{i}\left(1+\sum_{j=1}^{M}{t_{ij}r_{j}}+\frac{1}{2}\sum_{j=1}^{M}\sum_{k=1}^{M}t_{ij}t_{ik}r_{j}r_{k}+o^{2}\right),$
where there exists interaction terms $r_{j}r_{k}(j\not=k)$ of different mixture proportions.
Figure 16: Data mixing laws can model the language modeling performance of mixing correlated domains with different proportions. We train 70M models on the mixtures of ”Wikipedia+ CommonCrawl” and ”Wikipedia+ArXiv” for 15k steps. We validate on the two domains separately and fit the relationship between mixture proportions and validation losses with Eqn. 6.
Empirically, we evaluate the effectiveness of our data mixing laws in modeling domain interactions by examining their ability to predict language modeling performance when mixing two correlated domains. Specifically, we construct two synthetic data domains with deliberate overlap. The first domain consists of 50% Wikipedia and 50% CommonCrawl data, while the other domain comprises 50% Wikipedia and 50% ArXiv content. In this case, increasing the proportion of one domain necessarily increases the shared Wikipedia component. Therefore, the contribution of a training domain on target losses is coupled with the proportion of the other domain given their joint contribution through Wikipedia. As demonstrated in Fig.16, our proposed mixing law (Eqn.6) successfully models the language modeling performance across various mixing ratios of these correlated domains.
F.3 Extra Validation on Scaling Laws Prediction
Figure 17: The scaling law of training steps accurately extrapolates to 6.25x more steps. We fit $L(S)=E+B/S^{\beta}$ with 40k training steps of a 1B model and validate the prediction on language modeling performance up to 250k steps.
We discuss the computation that our prediction method with nested scaling laws requires. In particular, the cost primarily depends on how much scaling laws can accurately extrapolate.
Specifically, we need to train $N$ different data mixtures on model sized $N_{1},N_{2},\dots,N_{K}$ for $S_{0}$ steps to predict the model performance of different data mixtures trained with a model with $N_{target}$ parameters for $S_{target}$ steps. The total extra computational overhead relative to direct training is $N\frac{S_{0}\sum_{i=1}^{K}N_{i}}{S_{target}N_{target}}$, where the fraction $\frac{S_{0}\sum_{i=1}^{K}N_{i}}{S_{target}N_{target}}$ represents computation saved through scaling law predictions. State-of-the-art scaling law prediction demonstrates that this fraction can be 1/100 to 1/1000 (OpenAI, 2023; Bi et al., 2024). Together with the typical value of $N$, which is 20 in our experiments, the overall method should require an extra 1/5 to 1/50 training computation expectedly.
Figure 18: The scaling law of model sizes accurately extrapolates to 70x larger models. We fit language modeling performance at convergence following (Kaplan et al., 2020) with $L(N)=B/N^{\alpha}+E$. The language modeling performance of 1.5B and 7.25B models are predicted with L(S).
Given that achieving accurate scaling law predictions remains a developing area, we would like to provide our preliminary investigation to support 100x to 1000x scaling. Fig. 17 shows the scaling prediction of training steps with the scaling law of training steps $L(S)$, where we fit with the first 40k steps and predict the model performance up to 250k steps. This shows that fitting with 40k steps accurately predicts the language modeling performance on 250k steps, which is 6.25x scaling. Additionally, Fig. 18 shows the scaling prediction of model sizes with L(N), where we fit with models smaller than 100M and find it accurately predicts model performance up to 7.25B, which is  72.5x scaling. Combining L(S) and L(N), we may achieve  450x scaling.
F.4 Comparison to Other Data Mixing Methods
We compare our method to representative data mixing methods, DoGE (Fan et al., 2024) and DoReMi (Xie et al., 2024a). As our experiment in Sec. 4.2, we train on RedPajama and validation on the Pile.
Figure 19: The evolution of mixture proportions when training the proxy model with the updating rule in the OOD setting of DoGE.
DoGE (Fan et al., 2024) contains a universal generalization setting, which assumes validating on the same data as training, and an OOD setting which targets any validation data. We experiment with both of them. For universal generalization, we refer to the data mixture provided by Fan et al. (2024). For the OOD setting, we follow the original paper to train a small proxy model (160M) for 10k steps and apply their online updating rule to adjust the data mixture, shown in Fig. 19. We also follow Fan et al. (2024) to calculate the average proportions along the training steps of the proxy model as the final optimized mixture.
For DoReMi (Xie et al., 2024a), which is only designed for general optimization without awareness of the validation data, we experiment on both its mixture proportion optimized with RedPajama and the Pile. For the mixture optimized with RedPajama, we adopt the result of DoReMi10k from Fan et al. (2024). For the mixture optimized on the Pile, we refer to the optimized Pile mixture in the original paper (Xie et al., 2024a) and adapt the mixture to the one for RedPajama according to the domain overlap.
Figure 20: Comparisons of the language modeling performance of different data mixtures. All models are 1B models trained for 100B tokens with the same hyperparameters and validated on the validation set of the Pile. Default: original data mixture from Touvron et al. (2023a). DoGE (Universal): DoGE for universal generalization, obtained from Fan et al. (2024). DoGE (OOD): OOD generalization setting of DoGE optimized with validation set of the Pile. DoReMi (RedPajama): DoReMi mixture optimized by training proxy model on RedPajama. DoReMi (Pile): DoReMi mixture optimized by training proxy model on the Pile and adapted for our training on RedPajama through the domain overlaps between two dataset. Specific proportions are in Fig. 21.
Specifically, for ArXiv, Wikipedia, Github, and StackExchange, we directly borrow the mixture proportion. CommonCrawl and C4 equally share the proportion of Pile-CC. The proportion of Books is obtained as the sum of Books3 and BookCorpus2 in the Pile. We renormalize the proportions of these domains to ensure they sum up to 1.
Fig. 21 summarizes the final mixture proportion we use for different setups. We train all models for 100B tokens at the model size 1B. The outcome performance is in Fig. 20 which shows the mixture provided by our data mixing law indeed archives the lowest validation loss.
Figure 21: Specific mixture proportions on Redpajama from different data mixture optimization methods.
Appendix G Loss Prediction Results with Nested Scaling Laws
Fig. 22 shows the prediction results of nested use of scaling laws in Sec. 4.2. The result demonstrates plausible reference on the relative scale of losses on both the overall validation data and different validation domains. The optimized mixtures perform better in most domains. While the overall loss helps optimize the overall performance, losses on different domains show model capabilities in various aspects. Our result indicates that, by tuning data mixtures, it is possible to improve specific model capabilities without sacrificing others, consistent with the findings of Xie et al. (2024a).